# Variational Diffusion Posterior Sampling with Midpoint Guidance

### v1

@@ C.2 Warm start
10:     $Z_{\ell_{k}},Z_{k}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{N}(\mathrm{0}_{d},\mathrm{I}_{d})$
11:     $\hat{X}_{\ell_{k}}\leftarrow\hat{\mu}_{\ell_{k}}+\mathrm{diag}(\mathrm{e}^{\hat{\rho}_{\ell_{k}}})Z_{\ell_{k}}$
@@
13:        $(X_{k},\hat{X}_{1})\leftarrow\mathsf{WarmStart}(k,\hat{X}_{\ell_{k}},X_{k+1},M_{k})$
14:        $\hat{X}_{0}\leftarrow\hat{X}_{1}$
@@
16:        $X_{k}\leftarrow m_{k|{\ell_{k}},k+1}(\hat{X}_{\ell_{k}},X_{k+1})+v^{1/2}_{k|{\ell_{k}},k+1}Z_{k}$
17:        $\hat{X}_{0}\leftarrow\smash{m^{\theta}_{0|{\ell_{k}}}}(X_{\ell_{k}})$
@@
20:  Output: $\hat{X}_{0}$
@@ D.3 Gaussian mixtures
In this section, we elaborate more on the Gaussian mixture experiment in Table 4, where we consider a linear inverse problem with a Gaussian mixture as prior. Therefore, the potential is $g:x\mapsto\mathcal{N}(y;\mathrm{A}x,\sigma_{y}^{2}\mathrm{I}_{d_{y}})$. Recall that a Gaussian mixture with $C\in\mathbb{N}$ components, whose weights, means, and covariances are $w_{i}>0$, $m_{i}\in\mathbb{R}^{{d}}$, and $\sigma^{2}_{i}\mathrm{I}_{{d}}$, respectively, has the density
$q(x)=\sum_{i=1}^{C}w_{i}\mathcal{N}(x;m_{i},\sigma^{2}_{i}\mathrm{I}_{{d}})\,.$
@@
$\smash{m_{0|k}}(x_{k})=\big{(}x_{k}+(1-\alpha_{k})\nabla\log q_{k}(x_{k})\big{)}/\sqrt{\alpha_{k}},$
@@
$q_{k}(x_{k})$ $=\int q(x_{0})q_{k|0}(x_{k}|x_{0})\ \mathrm{d}x_{0}=\sum_{i=1}^{C}w_{i}\int\mathcal{N}(x_{0};m_{i},\sigma^{2}_{i}\mathrm{I}_{{d}})q_{k|0}(x_{k}|x_{0})\ \mathrm{d}x_{0}$
$=\sum_{i=1}^{C}w_{i}\mathcal{N}(x_{k};\sqrt{\alpha_{k}}m_{i},(\alpha_{k}\sigma^{2}_{i}+v_{k})\mathrm{I}_{{d}})\,.$
@@
$\pi(x)\propto g(x)q(x)\propto\sum_{i=1}^{C}\bar{w}_{i}\mathcal{N}(x;\bar{m}_{i},\bar{\Sigma}_{i})\,,$
@@
$\bar{w}_{i}$ $=w_{i}\mathcal{N}(y;\mathrm{A}m_{i},\sigma_{y}^{2}\mathrm{I}_{{d}_{y}}+\sigma^{2}_{i}\mathrm{A}\mathrm{A}^{\top})\,,$
$\bar{\Sigma}_{i}$ $=\big{(}(1/\sigma^{2}_{i})\mathrm{I}_{{d}}+(1/\sigma_{y}^{2})\mathrm{A}^{\top}\mathrm{A}\big{)}^{-1}\,,$
$\bar{m}_{i}$ $=\bar{\Sigma}_{i}\big{(}(1/\sigma_{y}^{2})A^{\top}y+(1/\sigma^{2}_{i})m_{i}\big{)}\,,$
@@
We consider two setups where ${d}\in\{20,200\}$ and generate Gaussian mixtures of $C=25$ components with means $(m_{i})_{i=1}^{C}=\{(8i,8j,\ldots,8i,8j)\in\mathbb{R}^{{d}}:(i,j)\in\llbracket-2,2\rrbracket^{2}\}$, unit covariances, and weights $w_{i}$ drawn from a uniform distribution on $[0,1]$ then normalized to sum to $1$. The linear inverse problem $(y,\mathrm{A})$ is generated by first drawing the matrix $\mathrm{A}\in\mathbb{R}^{1\times{d}}$ with entries i.i.d. from $\mathrm{N}(0,1)$, then sampling $x^{*}$ from the prior $q$ and finally computing $y=\mathrm{A}x^{*}+\sigma_{y}\varepsilon$ with $\varepsilon\in\mathrm{N}(0,1)$ to get the observation. The standard deviation of the inverse problem is fixed to $\sigma_{y}=0.05$ across all problem instances.
@@ D.4 Extended image experiments
We extend, in Tables 7 and 8, the results in Table 2 by including MGPS with $n=50$ and $n=100$. In the setting $n=100$, the runtime of MGPS is twice lower than that of DPS, PGDM and comparable to that of DiffPir, DDNM and RedDiff, see Figure 6. It is also outperforming them on most of the tasks, especially the nonlinear ones, as is seen in the table below. In the setting $n=50$, MGPS has the lowest runtime among all the methods while maintaining competitive performance.
@@
Table 7: Mean LPIPS values for various linear and nonlinear imaging tasks on the FFHQ $256\times 256$ dataset.
@@
Table 8: Mean LPIPS values for various linear and nonlinear imaging tasks on the ImageNet $256\times 256$ dataset.

### v2

@@ C.2 Warm start
10:     $\mathbf{Z}_{\ell_{k}},\mathbf{Z}_{k}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{N}(\mathbf{0}_{d},\mathbf{I}_{d})$
11:     $\hat{\mathbf{X}}_{\ell_{k}}\leftarrow\hat{\boldsymbol{\mu}}_{\ell_{k}}+\mbox{diag}(\mathrm{e}^{\hat{\boldsymbol{\rho}}_{\ell_{k}}})\mathbf{Z}_{\ell_{k}}$
@@
13:        $(\mathbf{X}_{k},\hat{\mathbf{X}}_{1})\leftarrow\mathsf{WarmStart}(k,\hat{\mathbf{X}}_{\ell_{k}},\mathbf{X}_{k+1},M_{k})$
14:        $\hat{\mathbf{X}}_{0}\leftarrow\hat{\mathbf{X}}_{1}$
@@
16:        $\mathbf{X}_{k}\leftarrow\boldsymbol{m}_{k|{\ell_{k}},k+1}(\hat{\mathbf{X}}_{\ell_{k}},\mathbf{X}_{k+1})+v^{1/2}_{k|{\ell_{k}},k+1}\mathbf{Z}_{k}$
17:        $\hat{\mathbf{X}}_{0}\leftarrow\smash{\boldsymbol{m}^{\theta}_{0|{\ell_{k}}}}(\mathbf{X}_{\ell_{k}})$
@@
20:  Output: $\hat{\mathbf{X}}_{0}$
C.3 Difference with DPS
In this section we detail the differences between MGPS and the DPS algorithm proposed in Chung et al. (2023). While DPS cannot be seen as an instantition of our methodology, we highlight the main differences by deriving a specific case of MGPS that closely resembles DPS.
We assume that ${\ell_{k}}=k$ for all $k\in\llbracket 1,n-1\rrbracket$ and that $p({\boldsymbol{y}}|\boldsymbol{x})=\mathcal{N}({\boldsymbol{y}};\mathcal{A}(\boldsymbol{x}),\sigma_{\boldsymbol{y}}^{2}\mathbf{I}_{d_{\boldsymbol{y}}})$ where $\mathcal{A}:\mathbb{R}^{d}\to\mathbb{R}^{d_{\boldsymbol{y}}}$. We consider the same optimization procedure as in Algorithm 1 but we: (i) optimize only the mean parameter $\hat{\boldsymbol{\mu}}_{k}$ and fix the covariance of the variational approximation to $v_{k|k+1}\mathbf{I}_{d}$, (ii) perform a single stochastic optimization step.
Following that, the initialization strategy (3.9) boils down to setting
$\lambda^{{\boldsymbol{\varphi}}_{k}}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})=p^{\theta}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})$ (C.4)
which means that the inital mean parameter is $\hat{\boldsymbol{\mu}}^{0}_{k}\vcentcolon=\boldsymbol{m}_{k|0,k+1}(\smash{\boldsymbol{m}^{\theta}_{0|k+1}}(\boldsymbol{x}_{k+1}),\boldsymbol{x}_{k+1})$. With this initialization ${\boldsymbol{\varphi}}_{k}\mapsto\mathsf{KL}(\lambda^{{\boldsymbol{\varphi}}_{k}}_{k|k+1}(\cdot|\boldsymbol{x}_{k+1})\parallel p^{\theta}_{k|k+1}(\cdot|\boldsymbol{x}_{k+1}))$ has its global optimum attained at ${\boldsymbol{\varphi}}_{k}=\hat{\boldsymbol{\mu}}^{0}_{k}$. Thus, the gradient (C.1) writes $\nabla_{{\boldsymbol{\varphi}}_{k}}\mathcal{L}_{k}({\boldsymbol{\varphi}}_{k};\boldsymbol{x}_{k+1})_{|\hat{\boldsymbol{\mu}}^{0}_{k}}=-\mathbb{E}\big{[}\nabla_{{\boldsymbol{\varphi}}_{k}}\log\hat{p}^{\theta}_{k}({\boldsymbol{y}}|\hat{\boldsymbol{\mu}}_{k}+\sqrt{v_{k|k+1}}\mathbf{Z})_{|\hat{\boldsymbol{\mu}}^{0}_{k}}\big{]}.$ Next, updating the parameters of the variational approximation using a single stochastic gradient descent step with the gradient estimate $-\nabla_{{\boldsymbol{\varphi}}_{k}}\log\hat{p}^{\theta}_{k}({\boldsymbol{y}}|\hat{\boldsymbol{\mu}}_{k}+\sqrt{v_{k|k+1}}\mathbf{Z}_{k})_{|\hat{\boldsymbol{\mu}}^{0}_{k}}$, where $\mathbf{Z}_{k}\sim\mathrm{N}(\mathbf{0}_{d},\mathbf{I}_{d})$ and step-size
$\gamma_{k}=\frac{\zeta\sigma_{\boldsymbol{y}}^{2}}{\|{\boldsymbol{y}}-\mathcal{A}\big{(}\smash{\boldsymbol{m}^{\theta}_{0|k}}(\hat{\boldsymbol{\mu}}^{0}_{k}+\sqrt{v_{k|k+1}}\mathbf{Z}_{k})\big{)}\|}$
yields the variational approximation
$\lambda^{{\boldsymbol{\varphi}}^{*}_{k}}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})=\mathcal{N}(\boldsymbol{x}_{k};\hat{\boldsymbol{\mu}}^{0}_{k}-\zeta\nabla_{{\boldsymbol{\varphi}}_{k}}\|{\boldsymbol{y}}-\mathcal{A}\big{(}\smash{\boldsymbol{m}^{\theta}_{0|k}}(\hat{\boldsymbol{\mu}}_{k}+\sqrt{v_{k|k+1}}\mathbf{Z}_{k})\big{)}\|_{|\hat{\boldsymbol{\mu}}^{0}_{k}},v_{k|k+1}\mathbf{I}_{d})\,.$
In order to simply the expression we have used the fact that
$\frac{-\nabla_{{\boldsymbol{\varphi}}_{k}}\log\hat{p}^{\theta}_{k}({\boldsymbol{y}}|\hat{\boldsymbol{\mu}}_{k}+\sqrt{v_{k|k+1}}\mathbf{Z}_{k})_{|\hat{\boldsymbol{\mu}}^{0}_{k}}}{\|{\boldsymbol{y}}-\mathcal{A}\big{(}\smash{\boldsymbol{m}^{\theta}_{0|k}}(\hat{\boldsymbol{\mu}}^{0}_{k}+\sqrt{v_{k|k+1}}\mathbf{Z}_{k})\big{)}\|}=\frac{1}{\sigma_{\boldsymbol{y}}^{2}}\nabla_{{\boldsymbol{\varphi}}_{k}}\|{\boldsymbol{y}}-\mathcal{A}\big{(}\smash{\boldsymbol{m}^{\theta}_{0|k}}(\hat{\boldsymbol{\mu}}_{k}+\sqrt{v_{k|k+1}}\mathbf{Z}_{k})\big{)}\|_{|\hat{\boldsymbol{\mu}}^{0}_{k}}\,.$
Therefore, given $\mathbf{X}_{k+1}$, a draw from the variational approximation in MGPS is obtained following the update
$\mathbf{X}_{k}=\tilde{\mathbf{X}}_{k}-\zeta\nabla_{\boldsymbol{x}_{k}}\|{\boldsymbol{y}}-\mathcal{A}(\smash{\boldsymbol{m}^{\theta}_{0|k}}(\boldsymbol{x}_{k}))\|_{|\boldsymbol{x}_{k}=\tilde{\mathbf{X}}^{\prime}_{k}}\,,\quad(\tilde{\mathbf{X}}_{k},\tilde{\mathbf{X}}^{\prime}_{k})\overset{\mathrm{i.i.d.}}{\sim}p^{\theta}_{k|k+1}(\cdot|\mathbf{X}_{k+1})\,.$
On the other hand, the DPS transition is given by
$\lambda^{\mathsf{DPS}}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})\vcentcolon=\mathcal{N}(\boldsymbol{x}_{k};\hat{\boldsymbol{\mu}}^{0}_{k}-\zeta\nabla_{\boldsymbol{x}_{k+1}}\|{\boldsymbol{y}}-\mathcal{A}(\smash{\boldsymbol{m}^{\theta}_{0|k+1}}(\boldsymbol{x}_{k+1}))\|,v_{k|k+1}\mathbf{I}_{d})\,.$
and, hence a sample $\mathbf{X}^{\mathsf{DPS}}_{k}$ is drawn following
$\mathbf{X}^{\mathsf{DPS}}_{k}=\tilde{\mathbf{X}}_{k}-\zeta\nabla_{\boldsymbol{x}_{k+1}}\|{\boldsymbol{y}}-\mathcal{A}(\smash{\boldsymbol{m}^{\theta}_{0|k+1}}(\boldsymbol{x}_{k+1}))\|_{|\boldsymbol{x}_{k+1}=\mathbf{X}_{k+1}}\,,\quad\tilde{\mathbf{X}}_{k}\sim p^{\theta}_{k|k+1}(\cdot|\mathbf{X}_{k+1})\,.$
The difference between MGPS and DPS is in: (i) the diffusion step used for the denoiser ($k$ for MGPS, $k+1$ for DPS), (ii) the sample where the gradient is evaluated.
@@ D.3 Gaussian mixtures
In this section, we elaborate more on the Gaussian mixture experiment in Table 4, where we consider a linear inverse problem with a Gaussian mixture as prior. Therefore, the likelihood is $p({\boldsymbol{y}}|\cdot):\boldsymbol{x}\mapsto\mathcal{N}({\boldsymbol{y}};\mathbf{A}\boldsymbol{x},\sigma_{\boldsymbol{y}}^{2}\mathbf{I}_{d_{\boldsymbol{y}}})$. Recall that a Gaussian mixture with $C\in\mathbb{N}$ components, whose weights, means, and covariances are $w_{i}>0$, $\boldsymbol{m}_{i}\in\mathbb{R}^{{d}}$, and $\sigma^{2}_{i}\mathbf{I}_{{d}}$, respectively, has the density
$q(\boldsymbol{x})=\sum_{i=1}^{C}w_{i}\mathcal{N}(\boldsymbol{x};\boldsymbol{m}_{i},\sigma^{2}_{i}\mathbf{I}_{{d}})\,.$
@@
$\smash{\boldsymbol{m}_{0|k}}(\boldsymbol{x}_{k})=\big{(}\boldsymbol{x}_{k}+(1-\alpha_{k})\nabla\log q_{k}(\boldsymbol{x}_{k})\big{)}/\sqrt{\alpha_{k}},$
@@
$q_{k}(\boldsymbol{x}_{k})$ $=\int q(\boldsymbol{x}_{0})q_{k|0}(\boldsymbol{x}_{k}|\boldsymbol{x}_{0})\ \mathrm{d}\boldsymbol{x}_{0}=\sum_{i=1}^{C}w_{i}\int\mathcal{N}(\boldsymbol{x}_{0};\boldsymbol{m}_{i},\sigma^{2}_{i}\mathbf{I}_{{d}})q_{k|0}(\boldsymbol{x}_{k}|\boldsymbol{x}_{0})\ \mathrm{d}\boldsymbol{x}_{0}$
$=\sum_{i=1}^{C}w_{i}\mathcal{N}(\boldsymbol{x}_{k};\sqrt{\alpha_{k}}\boldsymbol{m}_{i},(\alpha_{k}\sigma^{2}_{i}+v_{k})\mathbf{I}_{{d}})\,.$
@@
$\pi(\boldsymbol{x})\propto g(\boldsymbol{x})q(\boldsymbol{x})\propto\sum_{i=1}^{C}\bar{w}_{i}\mathcal{N}(\boldsymbol{x};\bar{\boldsymbol{m}}_{i},\bar{\mathbf{\Sigma}}_{i})\,,$
@@
$\bar{w}_{i}$ $=w_{i}\mathcal{N}({\boldsymbol{y}};\mathbf{A}\boldsymbol{m}_{i},\sigma_{\boldsymbol{y}}^{2}\mathbf{I}_{{d}_{{\boldsymbol{y}}}}+\sigma^{2}_{i}\mathbf{A}\mathbf{A}^{\top})\,,$
$\bar{\mathbf{\Sigma}}_{i}$ $=\big{(}(1/\sigma^{2}_{i})\mathbf{I}_{{d}}+(1/\sigma_{\boldsymbol{y}}^{2})\mathbf{A}^{\top}\mathbf{A}\big{)}^{-1}\,,$
$\bar{\boldsymbol{m}}_{i}$ $=\bar{\mathbf{\Sigma}}_{i}\big{(}(1/\sigma_{\boldsymbol{y}}^{2})A^{\top}{\boldsymbol{y}}+(1/\sigma^{2}_{i})\boldsymbol{m}_{i}\big{)}\,,$
@@
We consider two setups where ${d}\in\{20,200\}$ and generate Gaussian mixtures of $C=25$ components with means $(\boldsymbol{m}_{i})_{i=1}^{C}=\{(8i,8j,\ldots,8i,8j)\in\mathbb{R}^{{d}}:(i,j)\in\llbracket-2,2\rrbracket^{2}\}$, unit covariances, and weights $w_{i}$ drawn from a uniform distribution on $[0,1]$ then normalized to sum to $1$. The linear inverse problem $({\boldsymbol{y}},\mathbf{A})$ is generated by first drawing the matrix $\mathbf{A}\in\mathbb{R}^{1\times{d}}$ with entries i.i.d. from $\mathrm{N}(0,1)$, then sampling $\boldsymbol{x}^{*}$ from the prior $q$ and finally computing ${\boldsymbol{y}}=\mathbf{A}\boldsymbol{x}^{*}+\sigma_{\boldsymbol{y}}\varepsilon$ with $\varepsilon\sim\mathrm{N}(0,1)$ to get the observation. The standard deviation of the inverse problem is fixed to $\sigma_{\boldsymbol{y}}=0.05$ across all problem instances.
@@ D.4 Extended image experiments
To strengthen our conclusions and calculate the FID, we reran MGPS and the two closest competitors on 1000 images for the following five tasks: half mask, JPEG, motion deblur, nonlinear deblur, and high dynamic range, using the three priors (FFHQ, ImageNet, FFHQ-LDM). The Section D.4 shows that MGPS significantly outperforms the other competitors, including for the FID, with no significant change in other metrics compared to what we calculated on a smaller dataset.
We extend, in Tables 8 and 9, the results in Table 2 by including MGPS with $n=50$ and $n=100$. In the setting $n=100$, the runtime of MGPS is twice lower than that of DPS, PGDM and comparable to that of DiffPir, DDNM and RedDiff, see Figure 7. It is also outperforming them on most of the tasks, especially the nonlinear ones, as is seen in the table below. In the setting $n=50$, MGPS has the lowest runtime among all the methods while maintaining competitive performance.
@@
Table 7: Mean and confidence intervals of LPIPS/PSNR/SSIM values and FID value for various linear and nonlinear imaging tasks on 1k images for the 3 priors. Best is in bold.
FFHQ ImageNet FFHQ-LDM
Metric MGPS DDNM DiffPir MGPS DDNM DiffPir MGPS ReSample
Half mask
FID 27.0 38.6 45.2 40.0 50.0 57.0 49.5 66.6
LPIPS $\mathbf{0.19\pm 0.00}$ $0.23\pm 0.00$ $0.25\pm 0.00$ $\mathbf{0.30\pm 0.00}$ $0.38\pm 0.01$ $0.40\pm 0.01$ $\mathbf{0.26\pm 0.00}$ $0.30\pm 0.00$
PNSR $15.9\pm 0.2$ $\mathbf{16.3\pm 0.2}$ $16.1\pm 0.3$ $15.0\pm 0.1$ $\mathbf{16.0\pm 0.1}$ $15.8\pm 0.1$ $\mathbf{15.6\pm 0.1}$ $15.7\pm 0.1$
SSIM $0.70\pm 0.00$ $\mathbf{0.74\pm 0.00}$ $0.72\pm 0.01$ $0.63\pm 0.00$ $\mathbf{0.68\pm 0.01}$ $0.67\pm 0.01$ $\mathbf{0.69\pm 0.00}$ $0.67\pm 0.00$
MGPS DDNM DiffPir MGPS DDNM DiffPir MGPS ReSample
Motion deblur
FID 29.7 36.7 77.0 35.3 55.0 87.3 44.6 51.8
LPIPS $\mathbf{0.12\pm 0.00}$ $0.17\pm 0.00$ $0.22\pm 0.00$ $\mathbf{0.20\pm 0.01}$ $0.40\pm 0.01$ $0.39\pm 0.01$ $\mathbf{0.19\pm 0.00}$ $0.20\pm 0.00$
PNSR $26.7\pm 0.1$ $24.1\pm 0.1$ $\mathbf{27.4\pm 0.1}$ $\mathbf{24.4\pm 0.1}$ $21.4\pm 0.1$ $24.2\pm 0.1$ $26.4\pm 0.1$ $\mathbf{26.7\pm 0.1}$
SSIM $\mathbf{0.77\pm 0.00}$ $0.70\pm 0.01$ $0.71\pm 0.00$ $\mathbf{0.67\pm 0.01}$ $0.55\pm 0.01$ $0.61\pm 0.00$ $\mathbf{0.76\pm 0.00}$ $0.72\pm 0.00$
JPEG (QF = 2)
FID 31.6 87.6 109 61.4 128.8 92.8 45.0 65.3
LPIPS $\mathbf{0.15\pm 0.00}$ $0.37\pm 0.00$ $0.33\pm 0.01$ $\mathbf{0.40\pm 0.01}$ $0.60\pm 0.01$ $0.61\pm 0.00$ $\mathbf{0.21\pm 0.00}$ $0.26\pm 0.01$
PNSR $\mathbf{25.2\pm 0.1}$ $19.0\pm 0.2$ $24.5\pm 0.1$ $\mathbf{22.2\pm 0.1}$ $16.7\pm 0.1$ $22.2\pm 0.1$ $24.6\pm 0.1$ $\mathbf{24.8\pm 0.1}$
SSIM $\mathbf{0.73\pm 0.01}$ $0.55\pm 0.02$ $0.70\pm 0.00$ $\mathbf{0.60\pm 0.01}$ $0.41\pm 0.02$ $0.60\pm 0.01$ $\mathbf{0.71\pm 0.00}$ $0.66\pm 0.01$
Nonlinear deblur
FID 50.8 164 88.4 113 272 112 $\mathbf{69.2}$ 71.5
LPIPS $\mathbf{0.23\pm 0.01}$ $0.51\pm 0.02$ $0.68\pm 0.01$ $\mathbf{0.43\pm 0.01}$ $0.83\pm 0.01$ $0.66\pm 0.01$ $\mathbf{0.26\pm 0.01}$ $0.32\pm 0.01$
PNSR $\mathbf{24.3\pm 0.2}$ $16.2\pm 0.5$ $21.9\pm 0.1$ $\mathbf{22.2\pm 0.2}$ $9.9\pm 0.4$ $20.7\pm 0.2$ $23.9\pm 0.1$ $\mathbf{24.2\pm 0.1}$
SSIM $\mathbf{0.70\pm 0.01}$ $0.45\pm 0.02$ $0.42\pm 0.01$ $\mathbf{0.58\pm 0.01}$ $0.41\pm 0.01$ $0.241\pm 0.01$ $\mathbf{0.69\pm 0.01}$ $0.67\pm 0.01$
High dynamic range
FID 20.9 153 47.5 20.2 316 35.7 44.2 38.7
LPIPS $\mathbf{0.08\pm 0.01}$ $0.40\pm 0.04$ $0.20\pm 0.01$ $\mathbf{0.11\pm 0.01}$ $0.83\pm 0.02$ $0.20\pm 0.01$ $0.14\pm 0.00$ $\mathbf{0.12\pm 0.00}$
PNSR $\mathbf{27.0\pm 0.1}$ $18.7\pm 0.2$ $21.7\pm 0.1$ $\mathbf{26.3\pm 0.2}$ $9.9\pm 0.2$ $21.9\pm 0.1$ $25.5\pm 0.1$ $\mathbf{26.0\pm 0.1}$
SSIM $\mathbf{0.83\pm 0.01}$ $0.55\pm 0.04$ $0.72\pm 0.01$ $\mathbf{0.83\pm 0.01}$ $0.23\pm 0.02$ $0.71\pm 0.01$ $0.80\pm 0.01$ $\mathbf{0.83\pm 0.01}$
Table 8: Mean LPIPS/PSNR/SSIM values for various linear and nonlinear imaging tasks on the FFHQ $256\times 256$ dataset. Best is in bold and second best is underlined.
@@
PSNR  $\uparrow$
SR (×$\times 4$) 27.83 27.79 27.79 28.24 23.34 29.52 27.17 27.25
SR (×$\times 16$) 20.45 20.34 20.22 20.67 17.65 22.43 20.75 21.91
Box inpainting 21.55 22.22 22.68 18.39 21.13 22.35 21.96 21.79
Half mask 15.10 15.32 15.54 14.82 16.03 16.16 15.17 16.21
Gaussian Deblur 25.09 25.19 25.89 24.20 13.36 26.69 25.89 26.72
Motion Deblur 26.07 26.64 26.48 24.24 -- -- -- 27.58
JPEG (QF = 2) 25.00 25.23 24.94 18.50 12.76 -- -- 24.42
Phase retrieval 24.20 26.60 27.25 14.87 -- -- -- 24.85
Nonlinear deblur 24.16 24.21 24.37 15.89 -- -- -- 21.97
High dynamic range 24.77 26.07 27.74 16.83 -- -- -- 21.25
SSIM  $\uparrow$
SR (×$\times 4$) 0.81 0.80 0.79 0.81 0.50 0.85 0.77 0.70
SR (×$\times 16$) 0.58 0.57 0.55 0.59 0.38 0.67 0.60 0.62
Box inpainting 0.80 0.81 0.82 0.76 0.70 0.83 0.80 0.70
Half mask 0.69 0.70 0.71 0.66 0.56 0.73 0.67 0.65
Gaussian Deblur 0.71 0.73 0.75 0.69 0.14 0.77 0.73 0.76
Motion Deblur 0.77 0.78 0.77 0.71 -- -- -- 0.71
JPEG (QF = 2) 0.75 0.74 0.73 0.51 12.76 -- -- 0.71
Phase retrieval 0.74 0.78 0.79 0.43 -- -- -- 0.61
Nonlinear deblur 0.70 0.70 0.69 0.46 -- -- -- 0.42
High dynamic range 0.79 0.81 0.86 0.48 -- -- -- 0.71
Table 9: Mean LPIPS/PSNR/SSIM values for various linear and nonlinear imaging tasks on the Imagenet $256\times 256$ dataset. Best is in bold and second best is underlined.
@@
PSNR  $\uparrow$
SR (×$\times 4$) 24.68 24.70 24.77 23.52 15.67 25.55 24.26 24.24
SR (×$\times 16$) 18.56 18.42 18.04 18.22 15.80 20.43 19.37 19.95
Box inpainting 17.52 17.95 18.25 14.34 17.35 20.08 19.77 18.90
Half mask 14.98 15.14 15.60 14.65 14.36 17.06 15.79 16.96
Gaussian Deblur 22.56 21.96 20.10 21.20 9.93 23.29 22.10 23.27
Motion Deblur 23.91 25.03 24.50 21.59 -- -- -- 24.43
JPEG (QF = 2) 21.96 22.17 22.44 16.11 5.29 -- -- 22.15
Phase retrieval 16.36 16.94 18.10 14.40 -- -- -- 15.78
Nonlinear deblur 21.89 22.23 22.36 8.49 -- -- -- 20.76
High dynamic range 23.93 25.64 27.04 9.32 -- -- -- 22.88
SSIM  $\uparrow$
SR (×$\times 4$) 0.65 0.66 0.66 0.60 0.23 0.70 0.63 0.61
SR (×$\times 16$) 0.41 0.38 0.35 0.40 0.24 0.50 0.46 0.47
Box inpainting 0.71 0.74 0.76 0.70 0.62 0.78 0.74 0.67
Half mask 0.61 0.63 0.65 0.55 0.53 0.69 0.63 0.62
Gaussian Deblur 0.56 0.53 0.44 0.51 0.07 0.60 0.51 0.57
Motion Deblur 0.64 0.69 0.65 0.55 -- -- -- 0.60
JPEG (QF = 2) 0.59 0.60 0.60 0.39 0.01 -- -- 0.59
Phase retrieval 0.37 0.40 0.43 0.29 -- -- -- 0.26
Nonlinear deblur 0.57 0.58 0.58 0.24 -- -- -- 0.41
High dynamic range 0.75 0.81 0.84 0.25 -- -- -- 0.73
Table 10: Mean LPIPS/PSNR/SSIM values for various linear and nonlinear imaging tasks on FFHQ $256\times 256$ dataset with LDM prior. Best is in bold and second best is underlined.
MGPS ReSample PSLD MGPS ReSample PSLD MGPS ReSample PSLD
Task LPIPS  $\downarrow$ PSNR  $\uparrow$ SSIM  $\uparrow$
SR (×$\times 4$) 0.11 0.20 0.22 28.46 26.08 25.53 0.83 0.69 0.70
SR (×$\times 16$) 0.30 0.36 0.35 20.64 21.09 21.42 0.57 0.56 0.63
Box inpainting 0.16 0.22 0.26 22.94 18.80 20.39 0.79 0.75 0.66
Half mask 0.25 0.30 0.31 15.11 14.59 14.75 0.69 0.67 0.61
Gaussian Deblur 0.16 0.15 0.35 27.57 27.44 19.95 0.79 0.75 0.47
Motion Deblur 0.18 0.19 0.41 26.49 26.85 18.14 0.77 0.72 0.39
JPEG (QF = 2) 0.20 0.26 -- 24.75 24.33 -- 0.72 0.67 --
Phase retrieval 0.34 0.41 -- 22.21 19.05 -- 0.62 0.47 --
Nonlinear deblur 0.26 0.30 -- 23.79 24.44 -- 0.70 0.68 --
High dynamic range 0.15 0.15 -- 25.17 25.42 -- 0.79 0.81 --